OpenStudy (nsh4267):

Find the following values for the given functions. f(x)=x+4, g(x)=x^2 a) (f+g)(4) b) (f-g)(4) c) (fg)(4) d) (f/g)(4) Explain this to me I dont get it

OpenStudy (uri):

All those mean are add the two functions, subtract the two, multiply the two, and divide the two. a) (x+4) + (x^2) b) (x+4) - (x^2) c) (x+4)(x^2) d) (x+4)/(x^2) Just plug in 4 for x in all four of those and simplify

OpenStudy (nsh4267):

is that it? I thought it was more difficult then that

OpenStudy (uri):

That's it. Like for the multiplication as an example: (4+4)*(4^2) (8)*(16) = 128 Just do the same thing with the other questions.

OpenStudy (nsh4267):

ok how about this one. f(x)=-x^2, g(x)=6x-9 (f/g)(2)

OpenStudy (nsh4267):

I got -1.333333333

OpenStudy (uri):

-(2)^2/(6*2-9) -4/3 Id keep it a fraction, but -1.3333 is right.

OpenStudy (nsh4267):

changed it back to a fraction and it worked

OpenStudy (uri):

Good job